OpenStudy (anonymous):

viktoria travels from the point (4,6) to (6,8) then she travels from (6,8) to (8,10). How far is she from the point of start?

OpenStudy (anonymous):

Hint: find the distance between the points (4,6) and (8,10)

OpenStudy (anonymous):

Can u?

OpenStudy (anonymous):

yes then?

OpenStudy (anonymous):

but that is not the answer

OpenStudy (anonymous):

anyone please...i am stuck ...gosh why viktoria traveled lol

OpenStudy (anonymous):

please explain your work :D

OpenStudy (agent47):

d1=sqrt((6+8)^2+(4+6)^2) d1=sqrt(14^2+10^2) d1=sqrt(196+100) s1=sqrt(296) =17.2

OpenStudy (agent47):

i meant d1 at the end, not s1. Now d2..

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

i guess @sauravshakya 's answer seems correct

OpenStudy (agent47):

d2=sqrt(18^2+14^2) = sqrt(324+196) = sqrt(520) = 22.8

OpenStudy (agent47):

Total distance: d1+d2 = 17.2+22.8=40

OpenStudy (agent47):

ohh wait nvm

OpenStudy (agent47):

Means did not read the entire question: How far is she from the point of start?

OpenStudy (agent47):

do: sqrt(16^2+12^2)

OpenStudy (anonymous):

@Agent47 you are wrong

OpenStudy (anonymous):

\[D= \sqrt{(10-6)^2+(8-4)^2}= \sqrt{(4)^2+(4)^2} = \sqrt {32}\]

OpenStudy (agent47):

sqrt(256+144)=sqrt(400)=20 <FINAL ANSWER

OpenStudy (anonymous):

you are right @erica.d

OpenStudy (anonymous):

is it @sauravshakya ?

OpenStudy (anonymous):

@saiberz thanks

OpenStudy (agent47):

i failed yet again

OpenStudy (anonymous):

Yep...... thats it. @erica.d

OpenStudy (agent47):

:/

OpenStudy (anonymous):

thank you guys :)

OpenStudy (anonymous):

WELCOME

OpenStudy (agent47):

y cldn't anyone mention that i was using wrong signs??????????????!!!!!!!